OpenStudy (anonymous):

Solve. 7x2 – 35x = 168

OpenStudy (aroub):

\[7x^2 -35x=168 \] \[7x^2-35x-168=0 \] \[7(x^2-5x-24)=0 \] \[7 (x-8) (x+3) =0\] \[x= -3 & x=8 \]

OpenStudy (anonymous):

Solve. 3x2 = –21x – 15